Solve the equation √x−5=x−5.75 . You will find only one solution
Answer by lwsshak3(11628)

Solve the equation √x−5=x−5.75 . You will find only one solution
***
√x−5=x−5.75
√x=x-.75
square both sides
x=x^2-1.5x+.75^2
x^2-2.5x+.75^2=0
(x-.25)(x-2.25)=0
x=0.25
or
x=2.25
check:
For x=0.25
√(.25)-5=.25-5.75
-4.5≠5.5
0.25 is extraneous root
..
For x=2.25
√(2.25)-5=2.25-5.75
-3.5=-3.5
2.25 check ok